# Robust Hypothesis Generation: LLM-Automated Language Bias for Inductive Logic Programming

### v1

@@ Abstract
Automating robust hypothesis generation in open environments is pivotal for AI cognition. We introduce a novel framework integrating a multi-agent system, powered by Large Language Models (LLMs), with Inductive Logic Programming (ILP). Our system’s LLM agents autonomously define a structured symbolic vocabulary (predicates) and relational templates , i.e., language bias directly from raw textual data. This automated symbolic grounding (the construction of the language bias), traditionally an expert-driven bottleneck for ILP, then guides the transformation of text into facts for an ILP solver, which inductively learns interpretable rules. This approach overcomes traditional ILP’s reliance on predefined symbolic structures and the noise-sensitivity of pure LLM methods. Extensive experiments in diverse, challenging scenarios validate superior performance, paving a new path for automated, explainable, and verifiable hypothesis generation.
1 Introduction
Hypothesis generation—the process of forming systematic explanations from fragmented observations and iteratively validating them—plays a central role in advancing artificial intelligence [10]. This "generate-select" paradigm is fundamental not only in scientific discovery but also in practical AI tasks, such as diagnosing the root cause of software defects from error logs [27]. The capacity to continuously produce and verify hypotheses is thus crucial for building robust AI systems, especially when deployed in open environments or high-stakes domains like medical diagnosis and financial decision-making [12; 22].
@@ 1 Introduction
To harness the semantic strengths of LLMs while retaining the rigorous, verifiable outputs of ILP, we introduce a collaborative reasoning framework, shown in Figure 1. This framework first employs a multi-agent LLM system to automate the generation of a structured language bias, particularly the predicate system, directly from raw text. Subsequently, this LLM-generated bias guides the transformation of large-scale textual data into symbolic facts. This structured knowledge then empowers an ILP engine to conduct a robust, constrained search, yielding a globally coherent and optimal set of rules. Unlike prior approaches that often evaluate under idealized data conditions (e.g., assuming zero label noise or drawing conclusions such as ’8 examples are generally sufficient to describe the pattern’[20]), we provide a thorough analysis across more challenging data scenarios, substantiating our method’s superior efficacy and robustness.
@@
2 Preliminaries
In First-Order Logic (FOL), predicates are used to describe objects or the relationships between objects and can be classified according to the number of arguments, such as unary or binary. For example, the unary predicate isRed(x) denotes "x is red," and the binary predicate parent(x, y) denotes "x is a parent of y." Instantiating the arguments of a predicate to constants (e.g., parent(Alice, Bob)) yields an atom; if this atom is considered to be true, it is called a fact. Typically, the known information in a domain consists of several facts, which are regarded as directly usable background knowledge.
@@ 2 Preliminaries
3 Related Work
3.1 Inductive Logic Programming
@@
3.2 Hypothesis Generation Based on LLMs
@@ 4 Methodology
Figure 1: An simple illusration of our method and the difference between our method and other LLM-based hypothesis generation methods.
4.1 Predicate System Construction
@@
4.2 Symbolic Knowledge Encoding
Following predicate system finalization (Section 4.1), our Translator agent transforms natural language samples into Prolog facts. It parses each sample, mapping textual features to the established predicates. This translation proceeds in batches, circumventing LLM context limitations and the need for simultaneous full-dataset access common in other approaches. To bolster stability, translation failures trigger a retry (max two attempts). This systematic, batch-oriented encoding bridges natural language semantics with formal logic, enabling scalable conversion to a symbolic representation for ILP.
@@ 4.2 Symbolic Knowledge Encoding
4.3 ILP Learning
Upon completion of the symbolic knowledge encoding (Section 4.2), the comprehensive set of structured Prolog facts, along with the LLM-generated predicate system (Section 4.1), is provided as input to an ILP solver. We employ MAXSYNTH [7], which is an advanced solver that applies the Minimum Description Length (MDL) principle to balance rule complexity with noise coverage. This enables it to find globally optimal or near-optimal rule sets even with imperfect data, achieving both search efficiency and robustness, particularly in the presence of label noise. To enhance stability, if ILP solving fails, the algorithm restarts from predicate system design, with up to two attempts.
@@ 4.3 ILP Learning
5 Experiment Setup
5.1 Datasets and Baselines
Baselines: We consider two LLM-based inductive reasoning algorithms as baselines: HypoGeniC and Iterative Hypothesis Refinement (IHR). HypoGeniC generates hypotheses in natural language form and iteratively improves them using a bank of counterexamples. Iterative Hypothesis Refinement further enhances this process by generating, selecting, and refining hypotheses, and is also capable of producing executable code as candidate rules (see Appendix for details).
5.2 Experimental Variables
Rule Num refers to the number of underlying logical rules guiding the generation of facts for each sample. For example, a Zendo dataset with two rules may require "a red object on the left" (rule 1) or "a green object adjacent to a blue object" (rule 2), with positive samples satisfying either rule. Our experiments include rule sets containing 1, 2, or 3 rules.
5.3 Implementation Details
6 Experiments and Results
6.1 Main Experiments
@@
6.1.1 Comparative Analysis of Methods
@@
6.1.2 Dataset Complexity Analysis
@@
6.1.3 Model Dependency Analysis
@@
6.1.4 Visualization of Results
Table 2 illustrates how the three methods approach hypothesis generation differently on the same Zendo example. IHR generates code to verify hypotheses but achieves only 60% accuracy, as it struggles with complex relational reasoning despite its structured JSON input processing. HypoGeniC processes natural language descriptions directly but incorrectly infers a rule about "green pieces" (55% accuracy), demonstrating how pure LLM approaches can form plausible but incorrect generalizations. Our method, using the same natural language input as HypoGeniC, combines LLM capabilities with symbolic ILP solving to correctly identify the position-based pattern, achieving significantly higher accuracy. This example highlights how our hybrid approach effectively overcomes the limitations of both pure neural and symbolic reasoning techniques.
Method Input Output Acc
IHR Structured example: { "pieces": { "p36_0": { "orientation": "lhs", "size": "small", "color": "green", "position": [4, 5] }, "p36_1": { "orientation": "strange", "size": "small", "color": "red", "position": [2, 4] }, "p36_2": { "orientation": "lhs", "size": "large", "color": "blue", "position": [4, 6] }, "p36_3": { "orientation": "strange", "size": "medium", "color": "blue", "position": [0, 3] } }, "contacts": [] } Code example: def fn(x): pieces = x["pieces"] # Check if any pieces are in contact if x["contacts"]: return "No" # Get y-coordinates of all pieces y_coords = [piece["position"][1] for piece_id, piece in pieces.items()] # Check if all y-coordinates are even or all are odd all_even = all(y % 2 == 0 for y in y_coords) all_odd = all(y % 2 == 1 for y in y_coords) if all_even or all_odd: return "Yes" else: return "No" 60.0
HypoGeniC Text example:
"World 36: piece p36_0 is lhs-oriented, small and green, located at (4,5); piece p36_1 is strange-oriented, small and red, located at (2,4); piece p36_2 is lhs-oriented, large and blue, located at (4,6); piece p36_3 is strange-oriented, medium and blue, located at (0,3). World 36 is Zendo"
Text example:
"The world satisfies the Zendo rule if it contains at least two green pieces."
55.0
Ours Same with HypoGeniC Logic program: zendo_world(A):- has_piece(A,C), contacts(C,B), blue(B). zendo_world(A):- has_piece(A,B), strange_oriented(B), large(B). 85.0
6.2 Ablation Studies on Data-level Variables
@@
6.3 Overall Analysis
@@ 7 Conclusion
This study proposes an innovative multi-agent collaborative reasoning framework for complex knowledge discovery tasks. By integrating LLMs with ILP solver, our framework enables automatic hypothesis generation and verification through multi-agent collaboration. Our approach effectively overcomes the reliance of traditional ILP on expert-defined language bias, achieving an end-to-end automated pipeline from unstructured textual data to verifiable logical hypotheses. Extensive experiments on diverse datasets and challenging scenarios demonstrate that our framework consistently outperforms existing baselines under various data conditions, exhibiting superior performance and robustness. This work not only extends the application of ILP to unstructured text domains, but also provides a new paradigm for building interpretable hybrid AI reasoning systems, laying a solid foundation for automated knowledge discovery.
@@
Future Plan.
Future research will extend this framework to broader real-world scenarios, particularly tasks requiring hypothesis generation from large-scale unstructured texts. We plan to explore automatic identification of valuable questions and explanatory hypotheses across domains—analyzing human behavioral patterns in social sciences or generating scientific hypotheses about drug interactions and catalytic pathways in natural sciences.
@@ Appendix A MAXSYNTH
The core idea of the MAXSYNTH algorithm is to introduce the Minimum Description Length (MDL) principle into the search and optimization process of Inductive Logic Programming (ILP). Specifically, MDL advocates selecting the model with the lowest "description cost" when balancing hypotheses and data: on one hand, it seeks to use more compact rule expressions (reducing the complexity of the hypothesis), while on the other hand, it aims to minimize misclassification of observed data (reducing the data residual). In MAXSYNTH, this principle is quantified by computing the total cost of "program size + number of false positives + number of false negatives" for candidate logic programs. The program with the lowest total cost is selected as the output hypothesis, ensuring a balance between model simplicity and good data coverage, even in high-noise environments.
Unlike previous ILP algorithms, MAXSYNTH does not strictly pursue "full coverage of positive examples and zero coverage of negative examples." Instead, it allows a certain degree of error in order to ensure that a suitable explanation can still be found in complex or high-noise data scenarios. Additionally, compared to traditional greedy rule learning or closed searches based on manually defined predicate libraries, MAXSYNTH features a more flexible predicate invention mechanism and the ability to learn recursive rules. It adopts an iterative "generation–combination–constraint" process, integrating an optimal solver based on MaxSAT to progressively eliminate candidate programs that do not satisfy the optimal MDL criterion. This ensures that the remaining programs are globally optimal or near-optimal in terms of both size and misclassification cost. The process is theoretically safeguarded by a "noise-tolerant" constraint that prevents the elimination of any hypothesis that could potentially be an MDL-optimal solution, thereby guaranteeing the correctness and completeness of the algorithm.
@@ B.2 ZENDO
    zendo3(A) :- piece(A,D), blue(D), coord1(D,B), piece(A,C), coord1(C,B), red(C).
    zendo3(A) :- piece(A,D), contact(D,C), rhs(D), size(C,B), large(B).
    zendo3(A) :- piece(A,B), upright(B), contact(B,D), blue(D), size(D,C), large(C).
A world is classified as "Zendo" if it satisfies any of the defined logical rules; otherwise, it is classified as "Not Zendo."

### v2

# Hypothesis Generation via LLM-Automated Language Bias for ILP

@@ Abstract
Inductive Logic Programming (ILP) is a principled approach for generalizing regularities from data and constructing hypotheses as interpretable logic programs. However, a key limitation is its reliance on expert-crafted language bias—the predicate inventory, types, and mode declarations that delimit the search space. We propose hypothesis generation via LLM-automated language bias: multi-agent LLMs design the bias from raw text and translate descriptions into typed facts, and a robust ILP solver induces rules under a global consistency objective. This approach reduces traditional ILP’s reliance on predefined symbolic structures and the noise sensitivity of LLM-only pipelines that directly generate hypotheses as text or code. Extensive experiments in diverse, challenging scenarios validate superior performance, providing a practical, explainable, and verifiable route to hypothesis generation.
Introduction
Hypothesis generation—the process of forming systematic explanations from fragmented observations and iteratively validating them—plays a central role in advancing artificial intelligence [10]. This ”generate-test” paradigm is fundamental not only in scientific discovery but also in practical AI tasks, such as diagnosing the root cause of software defects from error logs [27]. The capacity to continuously produce and verify hypotheses is thus crucial for building robust AI systems, especially when deployed in open environments or high-stakes domains like medical diagnosis and financial decision-making [12; 22].
@@ Introduction
To harness the semantic strengths of LLMs while retaining the rigorous, verifiable outputs of ILP, we introduce a collaborative reasoning framework, shown in Figure 1. This framework first employs a multi-agent LLM system to automate the generation of a structured language bias, particularly the predicate system, directly from raw text. Subsequently, this LLM-generated bias guides the transformation of large-scale textual data into symbolic facts. This structured knowledge then empowers an ILP engine to conduct a robust, constrained search, yielding a globally coherent and optimal set of rules. Unlike prior approaches that evaluate under idealized conditions (e.g., zero label noise or fixed few-shot claims such as “8 examples suffice” [20]), we conduct a controlled robustness study along four axes: (i) surface-template perturbations (paraphrases, clause/order shuffling, light numeric fuzzing), (ii) label noise with flip rates $p!\in!{0.05,0.10,0.20}$, (iii) class imbalance with positive:negative ratios up to $1!:!9$, and (iv) relational complexity (unary SHOES vs. relational ZENDO). We further assess model-agnostic stability by swapping LLM backends and reporting cross-model variance, substantiating our method’s superior efficacy and robustness.
@@
Preliminaries
In First-Order Logic (FOL), predicates are used to describe objects or the relationships between objects and can be classified according to the number of arguments, such as unary or binary. For example, the unary predicate isRed(x) denotes ”x is red,” and the binary predicate parent(x, y) denotes ”x is a parent of y.” Instantiating the arguments of a predicate to constants (e.g., parent(Alice, Bob)) yields an atom; if this atom is considered to be true, it is called a fact. Typically, the known information in a domain consists of several facts, which are regarded as directly usable background knowledge.
@@
Related Work
Inductive Logic Programming
@@
Hypothesis Generation Based on LLMs
@@
Figure 1: Illustration of our pipeline and example rule. The LLMs produce an ILP language bias and typed facts, and a verifiable ILP solver induces rules under a global-consistency objective. Example hypothesis returned by the solver: zendo(A) :- piece(A,C), size(C,B), blue(C), small(B), contact(C,D), red(D). Meaning (variable types): AA: scene, $C,D$: pieces, $B$: size attribute. The rule states that $zendo(A)$ holds iff there exists a blue piece $C$ in scene AA that is small (via $size(C,B)\,\wedge\,small(B)$) and is in contact with a red piece $D$.
Predicate System Construction
@@
Symbolic Knowledge Encoding
Following predicate system finalization (Section Predicate System Construction), our Translator agent transforms natural language samples into Prolog facts. It parses each sample, mapping textual features to the established predicates. This translation proceeds in batches, circumventing LLM context limitations and the need for simultaneous full-dataset access common in other approaches. To bolster stability, translation failures trigger a retry (max two attempts). This systematic, batch-oriented encoding bridges natural language semantics with formal logic, enabling scalable conversion to a symbolic representation for ILP.
@@
ILP Learning
Upon completion of the symbolic knowledge encoding (Section Symbolic Knowledge Encoding), the comprehensive set of structured Prolog facts, along with the LLM-generated predicate system (Section Predicate System Construction), is provided as input to an ILP solver. We employ MAXSYNTH [7], which is an advanced solver that applies the Minimum Description Length (MDL) principle to balance rule complexity with noise coverage. This enables it to find globally optimal or near-optimal rule sets even with imperfect data, achieving both search efficiency and robustness, particularly in the presence of label noise. To enhance stability, if ILP solving fails, the algorithm restarts from predicate system design, with up to two attempts.
@@
Experiment Setup
Datasets, Baselines and Metrics
Baselines: We consider two LLM-based inductive reasoning algorithms as baselines: HypoGeniC[28] and Iterative Hypothesis Refinement (IHR)[20]. HypoGeniC generates hypotheses in natural language form and iteratively improves them using a bank of counterexamples. Iterative Hypothesis Refinement further enhances this process by generating, selecting, and refining hypotheses, and is also capable of producing executable code as candidate rules (see Appendix for details).
Metrics. We assess the reasonableness of a hypothesis by treating it as a decision results, using prediction Accuracy and Macro-F1.For text-outputting methods (HypoGeniC), following its original protocol we present the textual hypothesis plus the new example to the same LLM to obtain prediction label.
Experimental Variables
Rule Num refers to the number of underlying logical rules guiding the generation of facts for each sample. For example, a Zendo dataset with two rules may require ”a red object on the left” (rule 1) or ”a green object adjacent to a blue object” (rule 2), with positive samples satisfying either rule. Our experiments include rule sets containing 1, 2, or 3 rules.
Implementation Details
Experiments and Results
Main Experiments
Comparative Analysis of Methods
@@
Dataset Complexity Analysis
@@
Model Dependency Analysis
@@
Visualization of Results
Table Ablation Studies on Data-level Variables illustrates how the three methods approach hypothesis generation differently on the same Zendo example. IHR generates code to verify hypotheses but achieves only 60% accuracy, as it struggles with complex relational reasoning despite its structured JSON input processing. HypoGeniC processes natural language descriptions directly but incorrectly infers a rule about ”green pieces” (55% accuracy), demonstrating how pure LLM approaches can form plausible but incorrect generalizations. Our method, using the same natural language input as HypoGeniC, combines LLM capabilities with symbolic ILP solving to correctly identify the position-based pattern, achieving significantly higher accuracy. This example highlights how our hybrid approach effectively overcomes the limitations of both pure neural and symbolic reasoning techniques.
Ablation Studies on Data-level Variables
@@
Method Input Output Acc
IHR Structured example: 1{ 2 "pieces": { 3 "p36_0": { 4 "orientation": "lhs", 5 "size": "small", 6 "color": "green", 7 "position": [4, 5] 8 }, 9 "p36_1": { 10 "orientation": "strange", 11 "size": "small", 12 "color": "red", 13 "position": [2, 4] 14 }, 15 "p36_2": { 16 "orientation": "lhs", 17 "size": "large", 18 "color": "blue", 19 "position": [4, 6] 20 }, 21 "p36_3": { 22 "orientation": "strange", 23 "size": "medium", 24 "color": "blue", 25 "position": [0, 3] 26 } 27 }, 28 "contacts": [] 29} Code example: 1def fn(x): 2 pieces = x["pieces"] 3 4 # Check if any pieces are in contact 5 if x["contacts"]: 6 return "No" 7 8 # Get y-coordinates of all pieces 9 y_coords = [piece["position"][1] for piece_id, piece in pieces.items()] 10 11 # Check if all y-coordinates are even or all are odd 12 all_even = all(y % 2 == 0 for y in y_coords) 13 all_odd = all(y % 2 == 1 for y in y_coords) 14 15 if all_even or all_odd: 16 return "Yes" 17 else: 18 return "No" 60.0
HypoGeniC Text example:
”World 36: piece p36_0 is lhs-oriented, small and green, located at (4,5); piece p36_1 is strange-oriented, small and red, located at (2,4); piece p36_2 is lhs-oriented, large and blue, located at (4,6); piece p36_3 is strange-oriented, medium and blue, located at (0,3). World 36 is Zendo”
Text example:
”The world satisfies the Zendo rule if it contains at least two green pieces.”
55.0
Ours Same with HypoGeniC Logic program: 1zendo_world(A):- has_piece(A,C), contacts(C,B), blue(B). 2zendo_world(A):- has_piece(A,B), strange_oriented(B), large(B). 85.0
Overall Analysis
@@ Conclusion
This study proposes an innovative multi-agent collaborative reasoning framework for automatic hypothesis generation and verification. Our approach effectively reduces the reliance of traditional ILP on expert-defined language bias, achieving an end-to-end automated pipeline from unstructured textual data to verifiable logical hypotheses. Extensive experiments on diverse datasets and challenging scenarios demonstrate that our framework consistently outperforms existing baselines under various data conditions, exhibiting superior performance and robustness. This work not only extends the application of ILP to unstructured text domains, but also provides a new paradigm for building interpretable hybrid AI reasoning systems, laying a solid foundation for automated knowledge discovery.
@@ Appendix A MAXSYNTH
The core idea of the MAXSYNTH algorithm is to introduce the Minimum Description Length (MDL) principle into the search and optimization process of Inductive Logic Programming (ILP). Specifically, MDL advocates selecting the model with the lowest ”description cost” when balancing hypotheses and data: on one hand, it seeks to use more compact rule expressions (reducing the complexity of the hypothesis), while on the other hand, it aims to minimize misclassification of observed data (reducing the data residual). In MAXSYNTH, this principle is quantified by computing the total cost of ”program size + number of false positives + number of false negatives” for candidate logic programs. The program with the lowest total cost is selected as the output hypothesis, ensuring a balance between model simplicity and good data coverage, even in high-noise environments.
Unlike previous ILP algorithms, MAXSYNTH does not strictly pursue ”full coverage of positive examples and zero coverage of negative examples.” Instead, it allows a certain degree of error in order to ensure that a suitable explanation can still be found in complex or high-noise data scenarios. Additionally, compared to traditional greedy rule learning or closed searches based on manually defined predicate libraries, MAXSYNTH features a more flexible predicate invention mechanism and the ability to learn recursive rules. It adopts an iterative ”generation–combination–constraint” process, integrating an optimal solver based on MaxSAT to progressively eliminate candidate programs that do not satisfy the optimal MDL criterion. This ensures that the remaining programs are globally optimal or near-optimal in terms of both size and misclassification cost. The process is theoretically safeguarded by a ”noise-tolerant” constraint that prevents the elimination of any hypothesis that could potentially be an MDL-optimal solution, thereby guaranteeing the correctness and completeness of the algorithm.
@@ Appendix B Datasets Details
zendo3(A) :- piece(A,D), contact(D,C), rhs(D), size(C,B), large(B). zendo3(A) :- piece(A,B), upright(B), contact(B,D), blue(D), size(D,C), large(C).
A world is classified as ”Zendo” if it satisfies any of the defined logical rules; otherwise, it is classified as ”Not Zendo.”